# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 3.3. Toward generalization
    (3.176) $V_{N+1}(\eta_{2},z_{1},\dots,z_{N})=V_{N}(\eta_{2},z_{1},\dots,z_{N-1})+(\theta\lambda_{N+1}-\lambda_{1})\left(e^{z_{N-1}}-1-z_{N-1}\right)+\lambda_{1}\left(e^{z_{N}}-1-z_{N}\right),$
    with
    (3.177) $\begin{aligned} V_{N}(\eta_{2},z_{1},\cdots,z_{N-1})&=\theta\lambda_{2}(e^{\eta_{2}}-1-\eta_{2})+\theta\sum_{i=1}^{N-2}\lambda_{i+2}(e^{z_{i}}-1-z_{i})+\lambda_{1}(e^{z_{N-1}}-1-z_{N-1}).\end{aligned}$
    The time derivative $\dot{V}_{N+1}$ of the Lyapunov function $V_{N+1}$ satisfies
    (3.178) $\dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})=-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl{(}z_{i-1}\bigr{)}-c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)},$
    with a control of the form (3.168). ∎
@@
Under Assumption H6, by applying the same control of the form (3.164), the $N+1$-species system (3.146) is globally asymptotically stabilizable for $\psi_{i}\neq 0$, with a Lyapunov function of the form
(3.180) $V_{G}(\eta,\psi)=V_{{N+1}}(\eta_{2},z_{1},\cdots,z_{N-1})\Big{|}_{\ \psi_{i}\neq 0\ }\;+\;\frac{\gamma_{1}}{\sigma_{1}}h(G_{1}(\psi_{1}))\;+\;\frac{\gamma_{2}}{\sigma_{2}}h(G_{2}(\psi_{2}))\;+\;\cdots+\;\frac{\gamma_{N+1}}{\sigma_{N+1}}h(G_{N+1}(\psi_{N+1})).$
@@
Under Assumption H6, there exists a feedback control law under which the general non-transitive competition system (3.17) is globally asymptotically stable. Furthermore, the feedback control constructed remains uniformly bounded. In particular, the same control $u$ in (3.164) and the Lyapunov function $V_{G}$ in (3.154) ensures the global asymptotic stability of the $N+1$-species system (3.146). Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{G}(\eta,\psi)$ within the set
@@
    (3.182) $\begin{aligned} \dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl{(}z_{i-1}\bigr{)}-c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\\ &+\sum_{i=1}^{N+1}A_{i}\,\left(\hat{\phi}_{i}-\phi_{i}(\eta_{i})\right);\end{aligned}$
@@
    (3.183) $V_{G}(\eta,\psi)\leq$ ${}-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl{(}z_{i-1}\bigr{)}-c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}$
    (3.184) $+\sum_{i=1}^{N+1}\left(A_{i}\,\left(\phi_{i}(\eta_{i})+\lambda_{i}\right)-\gamma_{i}\right)\left(e^{G_{i}}-1\right).$
@@
    (3.185) $V_{G}(\eta,\psi)\leq$ $\ -\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl{(}z_{i-1}\bigr{)}-c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}$
    (3.186) $\ +\sum_{i=1}^{N+1}\left(C{i}\,|\phi_{i}(\eta_{i})+\lambda_{i}|-\gamma_{i}\right)\left(e^{G_{i}}-1\right).$
    establishing the result. ∎
@@ 3.4.2. Modeling
[…]
@@
The boundary conditions then link the aquatic and adult stages. Newly hatched individuals ($a=0$) derive from eggs laid by adult females: $I(0,t)=\int_{0}^{A}\beta\bigl{(}a,m(t)\bigr{)}\,F(a,t)\,\mathrm{d}a,$ where $\beta(a,m)$ is the fecundity rate, potentially modulated by male availability $m(t)=\displaystyle\int_{0}^{A}\lambda(a)M(a,t)da$. Upon maturation, the aquatic cohort gives rise to adult females or males according to $F(0,t)=r\int_{0}^{A}w(a)\,I(a,t)\,\mathrm{d}a,\quad M(0,t)=(1-r)\int_{0}^{A}w(a)\,I(a,t)\,\mathrm{d}a$ with $A=\max\{\tau,A^{*}\}.$
@@
Since model (3.187) is a non-autonomous logistic model, to carry out its qualitative analysis we may likewise replace the time-dependent functions $K(t)$, $\Gamma(t)$, and $\gamma(t)$ (see Section 3.4.3) by their respective mean values $K^{*}$, $\Gamma^{*}$, and $\gamma^{*}$. In the dynamic case, the functions $K(t)$, $\Gamma(t)$, and $\gamma(t)$ are assumed to be continuous and bounded on the interval $(0,T),$ namely
(3.188) $K(t),\;\Gamma(t),\;\gamma(t)\in L^{\infty}(0,T).$
@@
The mortality $\mu_{i}$ and fertility $\beta$ functions satisfy hypotheses $(H_{1})$ and $(H_{2})$, and the function $f$ meets condition $(H_{3})$.
@@ 3.4.3. Stability analysis
(3.210) $\pi_{0,I}(a)=\displaystyle\int_{a}^{A}\beta(s,m)e^{\int_{s}^{a}(\zeta_{I}+\mu_{I}(l,p)dl}ds,$
@@
$+\left\langle\pi_{0,F}(a),(\zeta_{F}-\displaystyle\int_{0}^{A}\gamma(t)F(a,t)da)F(a,t)\right\rangle$
@@
$+\left\langle\pi_{0,M}(a),(\zeta_{M}-\displaystyle\int_{0}^{A}\gamma(t)M(a,t)da)M(a,t)\right\rangle$
@@
$\left\langle\pi_{0,M}(a),\partial_{t}M(a,t)-(\zeta_{M}-\displaystyle\int_{0}^{A}\gamma(t)M(a,t)da)M(a,t)\right\rangle=0$
@@
(3.219) $\begin{cases}\mathcal{D}^{*}\pi_{0,I}(a)=\partial_{a}\pi_{0,I}(a)-\pi_{0,I}(a)(\mu(a,p(t))+\zeta_{I})+r\pi_{0,F}(0)w(a)+(1-r)\pi_{0,M}(0)w(a),\quad\pi_{0,I}(A)=0,\\ \\ \mathcal{D}^{*}\pi_{0,F}(a)=\partial_{a}\pi_{0,F}(a)-\pi_{0,F}(a)(\mu(a)+\zeta_{F})+\pi_{0,I}(0)\beta(a,m),\qquad\pi_{0,F}(A)=0,\\ \\ \mathcal{D}^{*}\pi_{0,M}(a)=\partial_{a}\pi_{0,M}(a)-\pi_{0,M}(a)(\mu_{M}(a)+\zeta_{M}),\qquad\pi_{0,M}(A)=0.\end{cases}$
@@
(3.220) $\partial_{t}F(a,t)\;=\;\Bigl{(}\zeta_{F}-\!\!\int_{0}^{A}\gamma(t)\,F(a,t)\,da\Bigr{)}F(a,t),\;\quad\partial_{t}M(a,t)\;=\;\Bigl{(}\zeta_{M}-\!\!\int_{0}^{A}\gamma(t)\,M(a,t)\,da\Bigr{)}M(a,t)$
@@
(3.221) $\partial_{t}I(a,t)\;=\;\Bigl{(}\zeta_{I}-P(t)+\Gamma(t)-\tfrac{\Gamma(t)\gamma(t)}{K(t)}\!\!\int_{0}^{A}I(a,t)\,da\Bigr{)}I(a,t)$
@@
$\partial_{t}\eta_{I}(t)=\zeta_{I}-P(t)+\Gamma(t)-\dfrac{\Gamma(t)\gamma(t)}{K(t)}\displaystyle\int_{0}^{A}I(a,t)da=\zeta_{I}+R_{I}(t)$
@@
(3.222) $\partial_{t}\eta_{I}(t)=\zeta_{I}-P(t)+\Gamma(t)-\dfrac{\Gamma(t)\gamma(t)}{K(t)}e^{\eta_{I}}\displaystyle\int_{0}^{A}(1+\psi_{I}(t-a))I^{*}(a)da,$
@@
(3.223) $\psi_{I}(t)=\frac{I(0,t)\,e^{-\eta_{I}(t)}}{I^{*}(0)}\;-\;1\Longrightarrow\psi_{I}(t)=\displaystyle\int_{0}^{A}g_{F}(a)\psi_{F}(t-a)F^{*}(a)da$
@@
(3.224) $\begin{cases}\psi_{F}(t)=\displaystyle\int_{0}^{A}g_{I}(a)\psi_{I}(t-a)I^{*}(a)da,\\ \psi_{M}(t)=\displaystyle\int_{0}^{A}g_{I}(a)\psi_{I}(t-a)I^{*}(a)da,\end{cases}$
@@
For the stability analysis of Theorems 3.31 and 3.34, we consider the following assumptions:

### v2

@@ 3.3. Toward generalization
    (3.138) $\begin{aligned} V_{N}(\eta_{2},z^{N}_{1},\cdots,z^{N}_{N-1})&=\theta\lambda_{2}(e^{\eta_{2}}-1-\eta_{2})+\theta\sum_{i=1}^{N-2}\lambda_{i+2}(e^{z^{N}_{i}}-1-z^{N}_{i})+\lambda_{1}(e^{z^{N}_{N-1}}-1-z^{N}_{N-1}).\end{aligned}$
    The time derivative $\dot{V}_{N}$ of the Lyapunov function $V_{N}$ satisfies
    (3.139) $\begin{aligned} \dot{V}_{N}(\eta_{2},z^{N}_{1},\dots,z^{N}_{N-1})_{\mid\psi_{i}\equiv 0}={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl(z^{N}_{i-1}\bigr)-c_{N}\,\phi_{1}^{2}\bigl(z^{N}_{N-1}\bigr)\end{aligned},$
    with a control of the form (3.127).
@@
Under Assumption H6, by applying the same control of the form (3.127), the NN- species system (3.123) is globally asymptotically stabilizable for $\psi_{i}\neq 0$, with a Lyapunov function of the form
(3.141) $V_{G}(\eta,\psi)=V_{{N}}(\eta_{2},z^{N}_{1},\cdots,z^{N}_{N-1})\;+\;\frac{\gamma_{1}}{\sigma_{1}}h(G_{1}(\psi_{1}))\;+\;\frac{\gamma_{2}}{\sigma_{2}}h(G_{2}(\psi_{2}))\;+\;\cdots+\;\frac{\gamma_{N}}{\sigma_{N}}h(G_{N}(\psi_{N})).$
@@
Under Assumption H6, there exists a feedback control law under which the general non-transitive competition system (3.16) is globally asymptotically stable. Furthermore, the feedback control constructed remains uniformly bounded. In particular, the same control $u$ in (3.127) and the Lyapunov function $V_{G}$ in (3.141) ensures the global asymptotic stability of the NN-species system (3.123). Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{G}(\eta,\psi)$ within the set
@@
    (3.143) $\begin{aligned} \dot{V}_{N}(\eta_{2},z^{N}_{1},\dots,z^{N}_{N-1})_{\mid\psi_{i}\neq 0}={}&-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl(z^{N}_{i-1}\bigr)-c_{N}\,\phi_{1}^{2}\bigl(z^{N}_{N-1}\bigr)+\sum_{i=1}^{N}A_{i}\,\left(\hat{\phi}_{i}-\phi_{i}(\eta_{i})\right);\end{aligned}$
@@
    (3.144) $V_{G}(\eta,\psi)\leq$ ${}-\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl(z^{N}_{i-1}\bigr)-c_{N}\,\phi_{1}^{2}\bigl(z^{N}_{N-1}\bigr)+\sum_{i=1}^{N}\left(A_{i}\,\left(\phi_{i}(\eta_{i})+\lambda_{i}\right)-\gamma_{i}\right)\left(e^{G_{i}}-1\right).$
@@
    (3.145) $V_{G}(\eta,\psi)\leq$ $\ -\theta c_{1}\,\phi_{2}^{2}(\eta_{2})-\theta\sum_{i=2}^{N-1}\,c_{\,i}\,\phi_{i+1}^{2}\bigl(z^{N}_{i-1}\bigr)-c_{N}\,\phi_{1}^{2}\bigl(z^{N}_{N-1}\bigr)-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl(z^{N}_{N-1}\bigr)-c_{N+1}\phi_{1}^{2}\bigl(z^{N}_{N}\bigr)$
    (3.146) $\ +\sum_{i=1}^{N}\left(C{i}\,|\phi_{i}(\eta_{i})+\lambda_{i}|-\gamma_{i}\right)\left(e^{G_{i}}-1\right).$
    establishing the result.
Proposition 3.28.
There exists $c>0$ such that a connected component of the sublevel set
$\mathcal{L}_{c}:=\{(\eta,\psi)\mid V_{G}(\eta,\psi)\leq c\}$
is contained in $\mathcal{K}_{N}$.
  • Proof of Proposition 3.28 :
    We select an initial vector
    $(\eta^{0},\psi^{0}),\qquad\text{with }\eta^{0}=(\varepsilon,0,0\cdots,0),\;\;\psi^{0}=(0,0,0,\cdots,0),$
    where $\varepsilon\neq 0$ is a sufficiently small real number, chosen such that $0<\varepsilon<\min\Big\{\ln\tfrac{\gamma_{1}}{C_{1}\lambda_{1}},\;\ln\tfrac{\gamma_{2}}{C_{2}\lambda_{2}},\;\ln\tfrac{\gamma_{3}}{C_{3}\lambda_{3}},\cdots,\ln\frac{\gamma_{N}}{C_{N}\lambda_{N}}\Big\}.$ Since $\phi_{1}$ is continuous with $\phi_{1}(0)=0$, one can further choose $\varepsilon>0$ small enough to guarantee $|c_{No}\phi_{1}(\varepsilon)|<\tfrac{u^{*}}{2}.$ Therefore,
    $u^{*}+c_{No}\phi_{1}(\varepsilon)>\tfrac{u^{*}}{2}>0,$
    which shows that $(\eta^{0},\psi^{0})\in\mathcal{K}_{N}$. In particular, this implies $\mathcal{K}_{N}\neq\varnothing$. Next, by continuity we have
    $V_{G}(\eta^{0},\psi^{0})=V_{N}(0,\cdots,0,\varepsilon)+\sum_{i=1}^{N}\frac{\gamma_{i}}{\sigma_{i}}\,h(G_{i}(\psi_{i}^{0})).$
    Since $V_{N}(0,\cdots,0,\varepsilon)\rightarrow 0$ as $\varepsilon\rightarrow 0$, and $h(0)=0$ whenever $G_{i}(\psi_{i}^{0})=0$, it follows that $V_{G}(\eta^{0},\psi^{0})$ can be made arbitrarily small.
    Hence, for any given $c>0$, there exists $\varepsilon>0$ sufficiently small such that $V_{G}(\eta^{0},\psi^{0})\leq c.$ Consequently, $(\eta^{0},\psi^{0})\in\mathcal{L}_{c}\cap\mathcal{K}_{N},$ which proves that $\mathcal{L}_{c}\cap\mathcal{K}_{N}\neq\varnothing.$
@@ 3.4.2. Modeling
System (3.147) describes the dynamics of an Anopheles mosquito population by distinguishing three age‑structured cohorts $a$: the aquatic population $I(a,t)$, adult females $F(a,t)$, and adult males $M(a,t)$.
@@
The boundary conditions then link the aquatic and adult stages. Newly hatched individuals ($a=0$) derive from eggs laid by adult females: $I(0,t)=\int_{0}^{A}\beta\bigl(a,m(t)\bigr)\,F(a,t)\,\mathrm{d}a,$ where $\beta(a,m)$ is the fecundity rate, potentially modulated by male availability $m(t)=\displaystyle\int_{0}^{A}\lambda(a)M(a,t)da$. Upon maturation, the aquatic cohort gives rise to adult females or males according to $F(0,t)=r\displaystyle\int_{0}^{A}w(a)\,I(a,t)\,\mathrm{d}a,\quad M(0,t)=(1-r)\displaystyle\int_{0}^{A}w(a)\,I(a,t)\,\mathrm{d}a$ with $A=\max\{\tau,A^{*}\}.$
@@
Since model (3.147) is a non-autonomous logistic model, to carry out its qualitative analysis we may likewise replace the time-dependent functions $K(t)$, $\Gamma(t)$, and $\gamma(t)$ (see Section 3.4.3) by their respective mean values $K^{*}$, $\Gamma^{*}$, and $\gamma^{*}$. In the dynamic case, the functions $K(t)$, $\Gamma(t)$, and $\gamma(t)$ are assumed to be continuous and bounded on the interval $(0,T),$ namely assuming hypothesis
$(HH):\begin{cases}K(t),\;\Gamma(t),\;\gamma(t)\in L^{\infty}(0,T),\\ \\ K(t)\geq\epsilon>0,\;\Gamma(t)\geq 0,\gamma(t)\geq 0,&\text{ in }(0,T).\end{cases}$
Moreover, we adopt the following standing hypotheses (unless otherwise stated):
$\textbf{(H33)}\begin{cases}\mu_{i}(a),\;\mu(a,p)\geq 0\quad\text{a.e. on }(0,A),\\[3.00003pt] \mu_{i}\in L^{1}_{\rm loc}(0,A),\quad\displaystyle\int_{0}^{A}\mu_{i}(a)\,da=+\infty,\end{cases}\qquad\textbf{(H44)}\begin{cases}\beta(.,m),\;w(.)\in L^{\infty}(0,A),\\[3.00003pt] \beta(a,m),\;w(a)\geq 0\ \text{a.e. on }(0,A).\end{cases}$
@@
The mortality $\mu_{i},$ fertility $\beta$ functions and $w$ the age‑dependent emergence rate satisfy hypotheses $(H33)$ and $(H44)$, and the function $f$ meets condition $(H3)$.
@@ 3.4.3. Stability analysis
(3.169) $\pi_{0,I}(a)=\displaystyle\int_{a}^{A}\beta(s,m)e^{\int_{s}^{a}(\zeta_{I}+\mu_{I}(l,p)dl}ds.$
@@
    $+\left\langle\pi_{0,F}(a),(\zeta_{F}-\displaystyle\int_{0}^{A}\gamma(t)F(a,t)da)F(a,t)\right\rangle,$
@@
    $+\left\langle\pi_{0,M}(a),(\zeta_{M}-\displaystyle\int_{0}^{A}\gamma(t)M(a,t)da)M(a,t)\right\rangle,$
@@
    $\left\langle\pi_{0,M}(a),\partial_{t}M(a,t)-(\zeta_{M}-\displaystyle\int_{0}^{A}\gamma(t)M(a,t)da)M(a,t)\right\rangle=0,$
@@
    (3.178) $\begin{cases}\mathcal{D}^{*}\pi_{0,I}(a)=\partial_{a}\pi_{0,I}(a)-\pi_{0,I}(a)(\mu(a,p(t))+\zeta_{I})+r\pi_{0,F}(0)w(a)+(1-r)\pi_{0,M}(0)w(a),\qquad\pi_{0,I}(A)=0,\\ \\ \mathcal{D}^{*}\pi_{0,F}(a)=\partial_{a}\pi_{0,F}(a)-\pi_{0,F}(a)(\mu(a)+\zeta_{F})+\pi_{0,I}(0)\beta(a,m),\hskip 18.49988pt\pi_{0,F}(A)=0,\\ \\ \mathcal{D}^{*}\pi_{0,M}(a)=\partial_{a}\pi_{0,M}(a)-\pi_{0,M}(a)(\mu_{M}(a)+\zeta_{M}),\hskip 18.49988pt\pi_{0,M}(A)=0.\end{cases}$
@@
    (3.179) $\partial_{t}F(a,t)\;=\;\Bigl(\zeta_{F}-\!\!\int_{0}^{A}\gamma(t)\,F(a,t)\,da\Bigr)F(a,t),\;\qquad\partial_{t}M(a,t)\;=\;\Bigl(\zeta_{M}-\!\!\int_{0}^{A}\gamma(t)\,M(a,t)\,da\Bigr)M(a,t),$
@@
    (3.180) $\partial_{t}I(a,t)\;=\;\Bigl(\zeta_{I}-P(t)+\Gamma(t)-\tfrac{\Gamma(t)\gamma(t)}{K(t)}\!\!\int_{0}^{A}I(a,t)\,da\Bigr)I(a,t),$
@@
    $\partial_{t}\eta_{I}(t)=\zeta_{I}-P(t)+\Gamma(t)-\dfrac{\Gamma(t)\gamma(t)}{K(t)}\displaystyle\int_{0}^{A}I(a,t)da=\zeta_{I}+R_{I}(t),$
@@
    (3.181) $\partial_{t}\eta_{I}(t)=\zeta_{I}-P(t)+\Gamma(t)-\dfrac{\Gamma(t)\gamma(t)}{K(t)}e^{\eta_{I}}\displaystyle\int_{0}^{A}(1+\psi_{I}(t-a))I^{*}(a)da.$
@@
    (3.182) $\psi_{I}(t)=\frac{I(0,t)\,e^{-\eta_{I}(t)}}{I^{*}(0)}\;-\;1\Longrightarrow\psi_{I}(t)=\displaystyle\int_{0}^{A}g_{F}(a)\psi_{F}(t-a)F^{*}(a)da.$
@@
    (3.183) $\begin{cases}\psi_{F}(t)=\displaystyle\int_{0}^{A}g_{I}(a)\psi_{I}(t-a)I^{*}(a)da,\\ \psi_{M}(t)=\displaystyle\int_{0}^{A}g_{I}(a)\psi_{I}(t-a)I^{*}(a)da.\end{cases}$
@@
For the stability analysis of Theorems 3.35 and 3.38, we consider the following assumptions: